1. Aspirin is itself an acid. (Its full name is acetylsalicylic acid.). It is neutralised by sodium hydroxide in this reaction:

C9H8O4 (aq) + NaOH (aq) →C9H7O4Na (aq) + H2O (l)

Some aspirin was prepared in the lab. Through titration, it was found that 4.00 g of the aspirin were neutralised by 17.5 cm3 of 1M sodium hydroxide solution. How pure was the aspirin sample?
1
Expert's answer
2020-06-09T07:19:56-0400

Mass of pure aspirin:

m= V(NaOH)*C(NaOH)*Mr(C9H8O4)=0.0175*1*180=3.15 g

%purity = 3.15/4.00 *100%=78.75 %

Answer: 78.75%